OpenStudy (anonymous):

evaluate the following rational expression and tell me how you do it x+7/x-3 for x=8

OpenStudy (anonymous):

3

OpenStudy (anonymous):

\[\frac{x+7}{x-3} = \frac{8+7}{8-3} = \frac{15}{5} = 3\]

OpenStudy (anonymous):

Sheg did it correctly. You just plug the 8 in where x is.

OpenStudy (anonymous):

thnx ladybug

OpenStudy (anonymous):

:) You're welcome!